1.      What is the oxidation state of the underlined element in (a) P4, (b) Al2O3, (c) MnO4 -, (d) NaH, (e) H2O2; (f) Fe3O4?


1
Expert's answer
2021-11-11T15:03:01-0500

a) "P_4" phosphorus bound to the same atom so oxidation number is zero.

b) "Al_2O_3 (2Al+3\times -2=0)", so oxidation number of Al is +3

c) "MnO_4^- ( Mn +4\times -2=-1)," oxidation number of Mn=+7

d) "NaH ( 1+H=0)", oxidation number of H=-1 and that of Na=+1

e) "H_2O_2(2\times (+1)+2O=0)", oxidation of oxygen =-1

f) "Fe_3O_4(2Fe+3\times -2=0)", so oxidation number of Fe=+3

Also for FeO, Fe+-2=0, therefore oxidation of Fe =+2

Thus oxidation of Fe= +3 and +2